OpenStudy (anonymous):

Simplify. f^-1g^-1 a. f/g b. g/f c. 1/fg d. fg

OpenStudy (johnweldon1993):

well f^-1 would be written as \[\frac{ 1 }{ f }\] g^-1 would be written as \[\frac{ 1 }{ g }\] so together you have \[\frac{ 1 }{ f } * \frac{ 1 }{ g }\] which is?

OpenStudy (anonymous):

fg?

OpenStudy (johnweldon1993):

\[\frac{ 1 * 1 }{ f * g }\]

OpenStudy (jhannybean):

f and g are two separate functions i believe

OpenStudy (anonymous):

1/fg idk im so confused

OpenStudy (jhannybean):

what are you confused about?

OpenStudy (johnweldon1993):

1/fg would be what I get yes.....but @Jhannybean what'd you mean by that? lol

OpenStudy (anonymous):

i dont understand how f^-1 went to 1/f and how g^-1 went to 1/g

OpenStudy (johnweldon1993):

ahh okay.....this is what happens when you want to turn a negative power ^-1 into a positive power ^1 what you do...is you put the function under a fraction under 1 a.k.a \[f ^{-1} = \frac{ 1 }{ f }\] *because no one likes working with negative exponents....so we turn them positive by doing this

OpenStudy (anonymous):

oh ok i think i get it now ok so for example if i have y^-3 it would be 3/y ???

OpenStudy (johnweldon1993):

not quite.....the function would STILL be under 1.... \[y ^{-3} = \frac{ 1 }{ y^3 }\] the exponent follows the 'y'....but just becomes positive

OpenStudy (johnweldon1993):

hmmm....another way of looking at it....you know what a reciprocal is correct?

OpenStudy (anonymous):

yes

OpenStudy (johnweldon1993):

perfect.....you know it is just the fraction flipped upside down \[\frac{ 1 }{ 2 } reciprocal = \frac{ 2 }{ 1 }\] like that well...when you have a function...raised to a negative power....a.k.a your \[f^{-1}\] look at the reciprocal of 'f' ...instead of 'f' it would be \[\frac{ 1 }{ f }\] and just bring the exponent with the 'f' \[f^{-1} = \frac{ 1 }{ f^1 } = \frac{ 1 }{ f }\] better?

OpenStudy (anonymous):

a little better, yes

OpenStudy (johnweldon1993):

now let me test you \[f^{-5} = ?\]

OpenStudy (anonymous):

1/f^5

OpenStudy (johnweldon1993):

perfect! :)

OpenStudy (anonymous):

yay

OpenStudy (johnweldon1993):

so now back to the original question...do you see why the answer would be \[\frac{ 1 }{ fg }\]

OpenStudy (anonymous):

because f^-1 is 1/f and g^-1 is 1/g

OpenStudy (johnweldon1993):

right....and then when you multiply them you get the 1/fg :)

OpenStudy (anonymous):

thank u for your help :D

OpenStudy (johnweldon1993):

anytime! :)